OpenStudy (anonymous):

B is the midpoint of line AC. AC = 100 . AB = x + y and BC = 3x-2y. Find x and y

OpenStudy (jdoe0001):

if AC = 100, and B is the midpoint, how long are AB and BC?

OpenStudy (anonymous):

50

OpenStudy (jdoe0001):

so AB = BC and each is 50 since AB = BC, and we know that AB = x + y, and that BC = 3x-2y we can say that x + y = 3x - 2y # can you solve that for "x"? the answer will be in terms of "y"

OpenStudy (anonymous):

I tried and i got 3y = 2x

OpenStudy (jdoe0001):

yeap, and solving for "x" will be \(\bf 3y = 2x \implies \cfrac{3y}{2} = x\) so now, we know that AB or BC are 50 so let's use say, BC = 3x-2y 3x - 2y = 50 but we also know that \(\bf x = \cfrac{3y}{2}\) so let's replace "x" in 3x - 2y for that, since they're equal \(\bf 3x - 2y = 50 \implies 3\left(\cfrac{3y}{2}\right) - 2y= 50 \implies \cfrac{9y}{2}-2y=50\\ \quad \\ \textit{multiplying both sides by 2, to get rid of the denominator}\\ 2 \times\left(\cfrac{9y}{2}-2y\right)=2\times(50) \implies 9y -4y = 100\) and if you solve for "y" there, you'll see what "y" is :)

OpenStudy (jdoe0001):

so, once you get the value for "y" you can use that, and plug in it on either equation, AB or BC and solve for "x" :)

OpenStudy (anonymous):

thank you sm uch!